Determine the area of the parallelogram determined by u(1,0) and v(0,1)
1
Expert's answer
2020-05-11T10:26:21-0400

We have given that vectors, "\textbf{u} =(1,0)" and "\textbf{v} =(0,1)" , Since, the vector cross product of any two vector gives the area of parallelogram. For simplicity let represent the vector in as usual way i.e "\textbf{u}=\hat{i} +0\hat{j} \: \& \: \textbf{v}=0\hat{i} + \hat{j}" ,Now we can calculate the vector cross product in two way,let's focus on the first method,

"\textbf{u} \times \textbf{v} = ||\textbf{u}||\: ||\textbf{v}|| \sin(\theta) \: \hat{n}", where "\theta" is the angle between "\textbf{u \& v}" , "||\textbf{u}||,||\textbf{v}||" are the norm or magnitude of vector "\textbf{u \& v}" ( here, "||\textbf{u}||=||\textbf{v}||=\sqrt{1^2 + 0^2} = 1" ),and "\hat{n}" is the direction of the area vector "\textbf{u} \times \textbf{v}" i.e "\hat{n}" is perpendicular to the plane of "\textbf{u \& v}" . Now, we have to find the angle "\theta" .Since, dot product of "\textbf{u \& v}" is "\textbf{u}\cdot\textbf{v}=1\cdot0 + 0\cdot1=0 \implies \textbf{u} \bot\textbf{v} \implies \theta = \frac{\pi}{2}" radian ,thus on calculating the vector cross product of "\textbf{u \& v}" , we get "||\textbf{u} \times \textbf{v}||=1". Hence, we are done.